If two triangles are equiangular,then their corresponding sides are in proportion. 
 

Grade:9

1 Answers

Saurabh Koranglekar
5 years ago
Dear student

The statement made is true
It can be verified by number of ways,

One such way is

(Sin A)/a =​(Sin B)/b =(Sin C)/c …..........1

(Sin A)/d =(Sin B)/e =(Sin C)/f …......2

from 1,2 we can see that sides are in proportion

Regards